# In Search of a Hidden Curve

### v1

@@ 2.5. Fugue in four themes: New proofs of Witten-Kontsevich and the $\lambda_{g}$ formulas
Here $[\ell;\hat{1}\hat{j}]=\{2,3,\dots,\hat{j},\dots,\ell\}$, and for a subset $I\subset[\ell]$ we write
$n_{I}=(n_{i})_{i\in I}\qquad{\text{and}}\qquad\tau_{n_{I}}=\prod_{i\in I}\tau_{n_{i}}.$
@@
Choose $n_{[\ell]}$ so that $|n_{[\ell]}|=n_{1}+n_{2}+\cdots+n_{\ell}=3g-3+\ell$. The degree of the left-hand side of (2.32) is $13(2g-2+\ell)+$ 1. So we compare the coefficients of $t_{1}^{2n_{1}+2}\prod_{j\geq 2}t_{j}^{2n_{j}+1}$ in the recursion formula. The contribution from the left-hand side of (2.32) is
${\langle}\tau_{n_{[\ell]}}{\rangle}_{g,\ell}(2n_{1}+1)!!\prod_{j\geq 2}(2n_{j}-1)!!.$! .
@@
$\sum_{j\geq 2}{\langle}\tau_{n_{1}+n_{j}-1}\tau_{n_{[\ell;\hat{1}\hat{j}]}}{\rangle}_{g,\ell-1}(2n_{1}+2n_{j}-1)!!\prod_{k\neq 1,j}(2n_{k}-1)!!.$! .
@@
$-(-1)^{g}(-1)^{2g-3+\ell}\sum_{i<j}{\langle}\tau_{n_{i}+n_{j}-1}\tau_{n_{[\ell;\hat{i}\hat{j}]}}\lambda_{g}{\rangle}_{g,\ell-1}\binom{n_{i}+n_{j}}{n_{i}}\prod_{i\geq 1}n_{i}!.$
@@
$b_{g}={\langle}\tau_{2g-2}\lambda_{g}{\rangle}_{g,1}\qquad g\geq 1$
@@
2.6. Coda: From a tree counting to the Lambert curve
@@
(2.43) $Dy=\frac{y}{1-y},\qquad D=x\,\frac{d}{dx}.$
@@
Although there is a difference between linear equation (1.3) and nonlinear equation (2.43), we consider them to be of the same nature, because they both lead to the spectral curve.
In this Coda, we deduce Differential Equation 2.43 from a purely combinatorial nature of the tree counting, and solve it to identify the Lambert curve, without ever appealing to the analysis of the Lagrange theorem.
We learn from the excellent book of Lovász et al. [62] that the $(0,1)$ Hurwitz number of degree $d$, $H_{0}\big{(}(d)\big{)}=\frac{d^{d-1}}{d!}$ of (2.34), is the number of rooted trees on $d$ nodes, counted in the stack sense. It means the reciprocal of the order of the automorphism group of each tree is used as a weight in counting. Cayley’s Theorem says that the number of all node-labeled trees on $d$ nodes is $d^{d-2}$, which is of course a positive integer. Since $d$ nodes are labeled in $d!$ different ways, the ratio $\frac{d^{d-2}}{d!}$ is the “number” of unlabeled trees. Since it is not an integer for $d\geq 2$, this “counting” is not the count of elements of a set. We are counting objects in a category, and automorphisms are taken into account.
@@
  • The book [62] mentioned above also talks about the set-theoretical count of the number of trees. There is no exact formula for that number. What becomes an important question in set-theoretical count is the asymptotic behavior of the number.
@@
  • Bingo! Yes, you go it. This is why we are calculating the Laplace transform! In both Catalan and Hurwitz cases, the Laplace transform led to the topology of the moduli spaces ${\mathcal{M}}_{g,n}$ and ¯${\overline{\mathcal{M}}}_{g,n}$. The idea here is similar to the Ehrhart polynomials and the Weil conjecture.
@@
Suppose now we have two rooted trees. We can joint these two roots with a new edge. The result is a new tree, where there is no root any more. But we have a particular edge, which did not exist in any of the two original trees. Thus we obtain a based tree, i.e., a tree in which a particular edge is chosen and declared it to be the base of the tree.
Figure 2.2. Construction of a based tree from two rooted trees.
The process if reversible. When you have a tree, and an edge has a name “base” on it, then we can simply remove it. Being a tree, the process produces two disjoint trees. The two ends of the removed edge will become the roots of the two trees. A tree on $n$ nodes has $n-1$ edges. So we obtain a bijective counting formula
(2.44) ${\frac{1}{2}}\sum_{\begin{subarray}{c}a+b=n\\ a\geq 1,b\geq 1\end{subarray}}\frac{a^{a-1}}{a!}\cdot\frac{b^{b-1}}{b!}=(n-1)\,\frac{n^{n-2}}{n!},\qquad n\geq 2.$
@@
$D\left({\frac{1}{2}}y^{2}\right)=xD\left(\frac{y}{x}\right)\quad\Longleftrightarrow\quad yy^{\prime}=\frac{xy^{\prime}-y}{x}\quad\Longleftrightarrow\quad Dy=\frac{y}{1-y}.$
@@
$(1-y)\;\frac{y^{\prime}}{y}=\frac{1}{x}\quad\Longrightarrow\quad x=ye^{-y}.$
@@
3. Variations of spectral curves in Higgs bundles and opers
@@
3.1. Variation 1: Moduli spaces of Higgs bundles and character varieties
@@
(3.1) ${\Sigma\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$$\scriptstyle{i}$${{T^{*}C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${\;C.}$
@@
$\Sigma=\left\{(z,\lambda)\in T^{*}C\;\left|\;z\in C,\;\lambda\in T^{*}_{z}C,\;\det\big{(}\lambda Id-\phi(z)\big{)}=0\right\}\right..$
@@
Since vector bundles on curves of genus 00 and $1$ behave differently from the curves of general type $g=g(C)>1$, let us restrict ourselves to the latter case for now. We also restrict our attention to vector bundles $E$ of rank $n$ with the fixed trivial determinant $\det(E)\cong{\mathcal{O}}_{C}$. It is natural to restrict the endomorphism sheaf for such a bundle to be traceless, because of the same reason of $\exp:sl_{n}({\mathbb{C}})\longrightarrow SL_{n}({\mathbb{C}})$. We denote the sheaf of traceless endomorphism by ${\text{\rm{End}}}_{0}(E)$, which is a rank $n^{2}-1$ locally free module over ${\mathcal{O}}_{C}$.
@@
A concrete example better explains this effect. Since ${\text{\rm{deg}}}(K_{C})=2g-2$, the canonical bundle has a square root $K_{C}^{\frac{1}{2}}$. Actually, the $\pm$ sign choice redundancy exists here, so there are $\left|H^{0}\big{(}C,{\mathbb{Z}}\big{/}2{\mathbb{Z}}\big{)}\right|=2^{2g}$ different choices of square roots. Any one of them is called a spin structure of $C$, and it appears in Riemann’s work as a $\theta$-characteristic. Choose a spin structure $K_{C}^{\frac{1}{2}}$ on $C$ and define $E=K_{C}^{\frac{1}{2}}\oplus K_{C}^{-{\frac{1}{2}}}$. Clearly $E$ is not a stable vector bundle. Since ${\text{\rm{End}}}(E)=K_{C}\oplus{\mathcal{O}}_{C}\oplus{\mathcal{O}}_{C}\oplus K_{C}^{-1}$, a non-zero quadratic differential
@@
$u=\sqrt{-1}\begin{bmatrix}1&&\omega\\ \\ 0&&-1\end{bmatrix}:\left(\begin{matrix}K_{C}^{\frac{1}{2}}\\ \oplus\\ K_{C}^{-{\frac{1}{2}}}\end{matrix}\right)\longrightarrow\left(\begin{matrix}K_{C}^{\frac{1}{2}}\\ \oplus\\ K_{C}^{-{\frac{1}{2}}}\end{matrix}\right),\qquad\det u=1.$
@@
where $[\alpha,\beta]=\alpha\beta\alpha^{-1}\beta^{-1}$ is the multiplicative commutator.
Figure 3.1. A homotopy basis for $\pi_{1}(C)$ of a curve of genus $2$. $[\alpha_{1},\beta_{1}]\cdot[\alpha_{2},\beta_{2}]=1$.
@@
(3.10) $B:=\bigoplus_{\ell=2}^{n}H^{0}\big{(}V,K_{C}^{\otimes\ell}\big{)},\quad\dim_{\mathbb{C}}\left(\sum_{\ell=2}^{n}\big{(}\ell(2g-2)-(g-1)\big{)}\right)=(g-1)(n^{2}-1).$
@@
(3.13) $X_{-}:=\big{[}\sqrt{s_{i-1}}\delta_{i-1,j}\big{]}_{ij}=\begin{bmatrix}0&0&\cdots&0&0\\ \sqrt{s_{1}}&&&&0\\ 0&\sqrt{s_{2}}&&&0\\ \vdots&&\ddots&&\vdots\\ 0&0&\cdots&\sqrt{s_{n-1}}&0\end{bmatrix},\qquad X_{+}:=X_{-}^{t},$
$H:=[X_{+},X_{-}],$
where $s_{i}:=i(n-i)$. $H$ is a diagonal matrix whose $(i,i)$-entry is $s_{i}-s_{i-1}=n-2i+1$, with $s_{0}=s_{r}=0$. The Lie subalgebra generated by $X_{+},X_{-},H$ is isomorphic to $sl(2,{\mathbb{C}})$, and is called the principal TDS of $SL(r,{\mathbb{C}})$.
@@
(3.14) $E_{0}:=\left(K_{C}^{\frac{1}{2}}\right)^{\otimes H}=\bigoplus_{i=1}^{n}\left(K_{C}^{\frac{1}{2}}\right)^{\otimes(n-2i+1)}=\bigoplus_{\ell=0}^{n-1}K_{C}^{-\frac{n-1}{2}}\otimes K_{C}^{\otimes\ell},\quad(\ell=n-i)$
@@
3.2. Variation 2: Higher order linear differential operators on a curve $C$
@@
  • To find special coordinate system on the curve $C$ that allows the expression (3.16) globally stays the same; and
@@
Here, the unknowns are the coordinate change $w=w(z)$ as a function in $z$, and the function $g(w)$. It is natural to assume that the 00-th order term $q$ of the differential operator satisfies $q(z)dz^{2}=q(w)dw^{2}$, i.e., it is a quadratic differential $q\in K_{C}^{\otimes 2}$. With these considerations, we can simplify (3.17) and obtain
@@
$\Longleftrightarrow$ $\begin{cases}2g_{w}w^{\prime}=\frac{w^{\prime\prime}}{w^{\prime}}\\ g_{ww}+g_{w}^{2}=0\;\end{cases}\Longleftrightarrow\quad\begin{cases}g(w)^{\prime}={\frac{1}{2}}(\log w^{\prime})^{\prime}\\ g_{ww}+g_{w}^{2}=0\;,\end{cases}$
@@
is the the Schwarzian derivative. We learn from Gunning [47] that every complex structure of $C$ admits a projective coordinate system, and that the transition function of this coordinate system
(3.20) $w(z)=\frac{az+b}{cz+d},\qquad\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,{\mathbb{C}}),$
@@
The calculation of (LABEL:2nd_calculation) can be extended to an operator $P$ of order $n$. The comparison of the $(n-1)$-order terms in
@@
$\binom{n}{2}(w^{\prime})^{n-2}w^{\prime\prime}\left(\frac{d}{dw}\right)^{n-1}=n(w^{\prime})^{n}g_{w}\left(\frac{d}{dw}\right)^{n-1}\Longleftrightarrow\quad g(w)^{\prime}=\frac{n-1}{2}(\log w^{\prime})^{\prime}.$
@@
3.3. Variation 3: Spectral curves and opers

### v2

@@ 2.5. New proofs of Witten-Kontsevich and the $\lambda_{g}$ formulas
Here $\ell;\hat{1}\hat{j}]=\{2,3,\dots,\hat{j}$$\dots,\ell\}$, and for a subset $I\subset[\ell]$ we write
$n_{I}=(n_{i})_{i\in I}\hskip 23.49976pt{\text{and}}\hskip 23.49976pt\tau_{n_{I}}=\prod_{i\in I}\tau_{n_{i}}.$
@@
Choose $n_{[\ell]}$ so that $|n_{[\ell]}|=n_{1}+n_{2}+\cdots+n_{\ell}=3g-3+\ell$. The degree of the left-hand side of (2.32) is $3(2g-2+\ell)+1$. So we compare the coefficients of $t_{1}^{2n_{1}+2}\prod_{j\geq 2}t_{j}^{2n_{j}+1}$ in the recursion formula. The contribution from the left-hand side of (2.32) is
${\langle}\tau_{n_{[\ell]}}{\rangle}_{g,\ell}(2n_{1}+1)!!\prod_{j\geq 2}(2n_{j}-1)!!.$
@@
$\sum_{j\geq 2}{\langle}\tau_{n_{1}+n_{j}-1}\tau_{n_{[\ell;\hat{1}\hat{j}]}}{\rangle}_{g,\ell-1}(2n_{1}+2n_{j}-1)!!\prod_{k\neq 1,j}(2n_{k}-1)!!.$
@@
$(-1)^{g}(-1)^{2g-3+\ell}\sum_{i<j}{\langle}\tau_{n_{i}+n_{j}-1}\tau_{n_{[\ell;\hat{i}\hat{j}]}}\lambda_{g}{\rangle}_{g,\ell-1}\binom{n_{i}+n_{j}}{n_{i}}\prod_{i\geq 1}n_{i}!.$
@@
$b_{g}={\langle}\tau_{2g-2}\lambda_{g}{\rangle}_{g,1}\hskip 23.49976ptg\geq 1$
@@
2.6. From a tree counting to the Lambert curve
@@
(2.43) $Dy=\frac{y}{1-y},\hskip 23.49976ptD=x\,\frac{d}{dx}.$
@@
In this Subsection, we deduce Differential Equation 2.43 from a purely combinatorial nature of the tree counting, and solve it to identify the Lambert curve, without appealing to the analysis of the Lagrange theorem.
We learn from the excellent textbook of Lovász et al. [66] that the $(0,1)$ Hurwitz number of degree $d$, $H_{0}\big((d)\big)=\frac{d^{d-1}}{d!}$ of (2.34), is the number of rooted trees on $d$ nodes, counted in the stack sense. It means the reciprocal of the order of the automorphism group of each tree is used as a weight in counting. Cayley’s Theorem says that the number of all node-labeled trees on $d$ nodes is $d^{d-2}$, which is of course a positive integer. Since $d$ nodes are labeled in $d!$ different ways, the ratio $\frac{d^{d-2}}{d!}$ is the “number” of unlabeled trees. Since it is not an integer for $d\geq 2$, this “counting” is not the count of elements of a set. We are counting objects in a category, and automorphisms are taken into account.
@@
  • The textbook [66] mentioned above also talks about the set-theoretical count of the number of trees. There is no exact formula for that number. What becomes an important question in set-theoretical count is the asymptotic behavior of the number.
@@
  • Bingo! Yes, you got it. This is why we are calculating the Laplace transform! In both Catalan and Hurwitz cases, the Laplace transform leads to the topology of the moduli spaces ${\mathcal{M}}_{g,n}$ and ${\overline{\mathcal{M}}}_{g,n}$. The idea here is similar to the Ehrhart polynomials and the Weil conjecture.
@@
Suppose now we have two rooted trees. We can join these two roots with a new edge. The result is a new tree, where there is no root any more. But we have a particular edge, which did not exist in any of the two original trees. Thus we obtain a based tree, i.e., a tree in which a particular edge is chosen and declared it to be the base of the tree.
Figure $2.$ Construction of a based tree from two rooted trees.
The process if reversible. When you have a tree, and an edge has a name “base” on it, then we can simply remove it. Being a tree, the process produces two disjoint trees. The two ends of the removed edge will become the roots of the two trees. A tree on $n$ nodes has n1n-1 edges. So we obtain a bijective counting formula
(2.44) ${\frac{1}{2}}\sum_{\begin{subarray}{c}a+b=n\\ a\geq 1,b\geq 1\end{subarray}}\frac{a^{a-1}}{a!}\cdot\frac{b^{b-1}}{b!}=(n-1)\,\frac{n^{n-2}}{n!},\hskip 23.49976ptn\geq 2.$
@@
$D\left({\frac{1}{2}}y^{2}\right)=xD\left(\frac{y}{x}\right)\qquad\Longleftrightarrow\qquad yy^{\prime}=\frac{xy^{\prime}-y}{x}\qquad\Longleftrightarrow\qquad Dy=\frac{y}{1-y}.$
@@
$(1-y)\;\frac{y^{\prime}}{y}=\frac{1}{x}\qquad\Longrightarrow\qquad x=ye^{-y}.$
@@
3. Spectral curves in Higgs bundles and opers
@@
3.1. Moduli spaces of Higgs bundles and character varieties
@@
[…]
@@
$\Sigma=\left\{(z,\lambda)\in T^{*}C\;\left|\;z\in C,\;\lambda\in T^{*}_{z}C,\;\det\big(\lambda I-\phi(z)\big)=0\right\}\right.$.
@@
Since vector bundles on curves of genus 0 and $1$ behave differently from the curves of general type $g=g(C)>1$, let us restrict ourselves to the latter case for now. We also restrict our attention to vector bundles $E$ of rank $n$ with the fixed trivial determinant $\det(E)\cong{\mathcal{O}}_{C}$. It is natural to restrict the endomorphism sheaf for such a bundle to be traceless, because of the same reason of $\exp:sl_{n}({\mathbb{C}})\longrightarrow SL_{n}({\mathbb{C}})$. We denote the sheaf of traceless endomorphism by ${\text{\rm{End}}}_{0}(E)$, which is a rank $n^{2}-1$ locally free module over ${\mathcal{O}}_{C}$.
@@
A concrete example better explains this effect. Since ${\text{\rm{deg}}}(K_{C})=2g-2$, the canonical bundle has a square root $K_{C}^{\frac{1}{2}}$. Actually, the $\pm$ sign choice redundancy exists here, so there are $\left|H^{0}\big(C,{\mathbb{Z}}\big/2{\mathbb{Z}}\big)\right|=2^{2g}$ different choices of the square root. Any one of them is called a spin structure of $C$, and it appears in Riemann’s work as a $\theta$-characteristic. Choose a spin structure $K_{C}^{\frac{1}{2}}$ on $C$ and define $E=K_{C}^{\frac{1}{2}}\oplus K_{C}^{-{\frac{1}{2}}}$. Clearly $E$ is not a stable vector bundle. Since ${\text{\rm{End}}}(E)=K_{C}\oplus{\mathcal{O}}_{C}\oplus{\mathcal{O}}_{C}\oplus K_{C}^{-1}$, a non-zero quadratic differential
@@
$u=\sqrt{-1}\begin{bmatrix}1&&\omega\\ \\ 0&&-1\end{bmatrix}:\left(\begin{matrix}K_{C}^{\frac{1}{2}}\\ \oplus\\ K_{C}^{-{\frac{1}{2}}}\end{matrix}\right)\longrightarrow\left(\begin{matrix}K_{C}^{\frac{1}{2}}\\ \oplus\\ K_{C}^{-{\frac{1}{2}}}\end{matrix}\right),\hskip 23.49976pt\det u=1.$
@@
where $\alpha,\beta]=\alpha\beta\alpha^{-1}\beta^{-1}$ is the multiplicative commutator.
Figure 3.1. A homotopy basis for $\pi_{1}(C)$ of a curve of genus $2$. $\alpha_{1},\beta_{1}]\cdot[\alpha_{2},\beta_{2}]=1$.
@@
(3.10) $B:=\bigoplus_{\ell=2}^{n}H^{0}\big(V,K_{C}^{\otimes\ell}\big),\qquad\dim_{\mathbb{C}}\left(\sum_{\ell=2}^{n}\big(\ell(2g-2)-(g-1)\big)\right)=(g-1)(n^{2}-1).$
@@
(3.13) $X_{-}=\big[\sqrt{s_{i-1}}\delta_{i-1,j}\big]_{ij}=\begin{bmatrix}0&0&\cdots&0&0\\ \sqrt{s_{1}}&&&&0\\ 0&\sqrt{s_{2}}&&&0\\ \vdots&&\ddots&&\vdots\\ 0&0&\cdots&\sqrt{s_{n-1}}&0\end{bmatrix},\qquad X_{+}=X_{-}^{t},$
$H=[X_{+},X_{-}],$
where $s_{i}:=i(n-i)$. $H$ is a diagonal matrix whose $(i,i)$-entry is $s_{i}-s_{i-1}=n-2i+1$, with $s_{0}=s_{r}=0$. The Lie subalgebra generated by $X_{+},X_{-},H$ is isomorphic to $sl(2,{\mathbb{C}})$, and is called a principal TDS of $SL(r,{\mathbb{C}})$.
@@
(3.14) $E_{0}:=\left(K_{C}^{\frac{1}{2}}\right)^{\otimes H}=\bigoplus_{i=1}^{n}\left(K_{C}^{\frac{1}{2}}\right)^{\otimes(n-2i+1)}=\bigoplus_{\ell=0}^{n-1}K_{C}^{-\frac{n-1}{2}}\otimes K_{C}^{\otimes\ell},\qquad(\ell=n-i)$
@@
3.2. Higher order linear differential operators on a curve $C$
@@
  • To find a coordinate system on the curve $C$ that allows the expression (3.16) stays globally the same; and
@@
Here, the unknowns are the coordinate change $w=w(z)$ as a function in $z$, and the function $g(w)$. It is natural to assume that the 0-th order term $q$ of the differential operator satisfies $q(z)dz^{2}=q(w)dw^{2}$, i.e., it is a quadratic differential $q\in K_{C}^{\otimes 2}$. With these considerations, we can simplify (3.17) and obtain
@@
[…]
@@
is the Schwarzian derivative. We learn from Gunning [49] that every complex structure of $C$ admits a projective coordinate system, and that the transition function of this coordinate system
(3.20) $w(z)=\frac{az+b}{cz+d},\hskip 23.49976pt\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in SL(2,{\mathbb{C}}),$
@@
The calculation of (LABEL:2nd_calculation) can be extended to an operator $P$ of order $n$. The comparison of the (n1)(n-1)-order terms in
@@
$\binom{n}{2}(w^{\prime})^{n-2}w^{\prime\prime}\left(\frac{d}{dw}\right)^{n-1}=n(w^{\prime})^{n}g_{w}\left(\frac{d}{dw}\right)^{n-1}\Longleftrightarrow\qquad g(w)^{\prime}=\frac{n-1}{2}(\log w^{\prime})^{\prime}.$
@@
3.3. Spectral curves and opers